[6] Polar coordinates of a point are given. Find the rectangular coordinates of each point. Note: If e is a special angle,
百 6 6 6
Answer #1

Solution- 0 (11) is equivalent to (11,-) 8-11, 05- A xt 0 coses 11 cos(-4)=0 Yarsino = 11 sin(-4) = 11*(-1) -11,8- 110,-11)(-19, -79) is equivalent to (14,4 - 3 205 rcoso = 14 (05 (-34) -1441-16) 2-712 > X xt 14 AE -3 5 y=-789 14 (2 2 √g 14 JZ 2 گ